Fidelity records the number of X-Trails sold each day. This data is used in calculating the following probability distribution of daily sales:
x P(x)
0 0.1
1 0.1
2 0.2
3 0.2
4 0.3
5 0.1
1) Find the cumulative distribution function (CDF)
Some help would be greatly appreciated, thanks!

Answer by Edwin McCravy(20063)

Fidelity records the number of X-Trails sold each day. This data is used in calculating the following probability distribution of daily sales:
1) Find the cumulative distribution function (CDF)
Some help would be greatly appreciated, thanks!
The word "cumulative" reminds you of 'accumulating', so you 'accumulate'
the P(x)'s as you go, by adding all the previous values to each one.

                x       P(x)   CDP(x)
                0       0.1    0.1 = 0.1                
                1       0.1    0.1+0.1 = 0.2
                2       0.2    0.1+0.1+0.2 = 0.4
                3       0.2    0.1+0.1+0.2+0.2 = 0.6
                4       0.3    0.1+0.1+0.2+0.2+0.3 = 0.9
                5       0.1    0.1+0.1+0.2+0.2+0.3+0.1 = 1.0

Edwin